# Selective Mixup for Debiasing Question Selection in Computerized Adaptive Testing

### v1

@@ 5.2. Performance in OOD Settings
Motivation. In most existing data-driven CAT studies, the meta set used to evaluate question selection performance is typically sampled directly from the examinee’s original interaction log. This procedure implicitly assumes that the observed response distribution accurately represents the examinee’s true question exposure or proficiency profile. However, in practice, this assumption is problematic. Since CAT adaptively selects questions based on estimated ability, examinees are only exposed to a small, biased subset of the entire question pool. Consequently, the meta set may reflect these selection-induced patterns rather than the examinee’s actual proficiency, leading to a biased and unrepresentative evaluation.
To better evaluate the effectiveness of our debiasing strategy in mitigating selection bias, we design an Out-of-Distribution (OOD) test setting. Unlike the conventional IID setup where the meta set is randomly sampled from the same distribution as the support set, our OOD protocol deliberately alters this distributional assumption. Specifically, for each examinee, we construct an unbiased meta set by enforcing a balanced ratio of correct and incorrect responses (1:1). The remaining interactions are then assigned to the support set, which naturally leads to a distribution shift between the two sets. Examinees whose data cannot satisfy this balancing constraint are excluded. In this way, we preserve the support–meta framework while ensuring that the evaluation truly reflects performance under distributional shifts. This controlled design provides a more rigorous and interpretable benchmark, making the OOD results highly indicative of a model’s robustness and generalization ability.
Our proposed method demonstrates strong performance across different settings, consistently surpassing baseline approaches on the Worst metric and outperforming most baselines on the Avg. metric across two distinct Cognitive Diagnostic Modules and two widely used public datasets. Specifically, on the ASSIST0910 dataset, our method achieves substantial relative improvements ranging from 2.72% to 21.5% on the Worst metric, and from 0.20% to 1.64% on the Avg. metric. Similarly, on the NIPS-EDU dataset, we observe performance gains of 1.16% to 32.69% on the Worst metric, and 0.11% to 0.21% on the Avg. metric, compared to the state-of-the-art (SOTA) results. Notably, the larger improvement margins under the IRT setting, as opposed to the NCDM setting, suggest that our method better complements the limitations of traditional models like IRT, thereby further highlighting the strengths of neural Cognitive Diagnostic Modules such as NCDM in mitigating individual performance variance. These results suggest that our approach not only provides reliable gains over ERM but also outperforms existing debiasing strategies, especially in the Worst metric, which is critical for robustness in adaptive testing.
@@ 5.3. Performance in IID Settings
As shown in Table 2 and Table 3, debiasing inevitably introduces a trade-off between IID and OOD performance. While methods such as Reweight and our approach are effective in mitigating bias and improving OOD generalization, they cannot surpass ERM or GroupDRO under IID settings, which are naturally more favorable to models optimized without debiasing constraints. Specifically, IRM achieves the best average performance on ASSIST2009 (0.6939), while ERM achieves the highest average on NIPS-EDU (0.6930). GroupDRO also maintains relatively strong IID results across both datasets. In contrast, our method focuses on improving robustness: it substantially outperforms the Reweight baseline in both Worst and Avg. accuracy under OOD conditions, while remaining competitive in-distribution. This confirms that debiasing methods, including ours, sacrifice a small amount of IID accuracy in exchange for much stronger OOD generalization, underscoring the trade-off inherent in fairness- and robustness-oriented learning.
@@ 5.4. Hyperparameter Analysis
We conduct a hyperparameter analysis of the mixup regularization weight $\omega$ on the NIPS-EDU dataset, evaluating its impact on Worst accuracy across two Cognitive Diagnosis Models and two CAT frameworks. As shown in Figure 4, most CDM-CAT combinations achieve their best performance when $\omega$ is set to 0.8 or 1.0. The overall trend follows a rise-then-fall pattern: performance initially improves as $\omega$ increases, peaking around 0.8 or 1.0, and then gradually declines beyond that point.Notably, our method demonstrates relatively stable performance across a wide range of $\omega$ values. This suggests that the model is not overly sensitive to the specific setting of the regularization strength, which enhances its practical applicability. Even when $\omega$ deviates slightly from the optimal value, the model maintains strong worst-group accuracy, indicating robust debiasing behavior and consistent generalization across configurations.
Figure 4. The effect of the regularization weight $\omega$ on Worst accuracy shows that performance peaks around 0.2–1.0 and remains relatively stable across a wide range.
5.5. Ablation Study
We conduct ablation studies to examine three mixup strategies: Mixup-Self, Mixup-Inner, and Mixup-B. In Mixup-Self, interactions are interpolated within the same examinee, which introduces synthetic variation into an individual learner’s profile and can occasionally improve robustness. Mixup-Inner interpolates the most similar examinees within the same attribute, thereby smoothing local representations and reducing overfitting to individual-specific patterns. In contrast, Mixup-B interpolates biased examinees (e.g., from Attribute A or C) with balanced counterparts (Attribute B). By explicitly synthesizing bias-conflicting interactions, this strategy enriches underrepresented patterns and more effectively mitigates selection bias. As shown in Figure 5, Mixup-B consistently achieves superior performance, particularly on the large-scale ASSIST0910 dataset, where the greater number of questions provides higher diversity and amplifies the benefits of cross-attribute mixing. These results highlight both the effectiveness of cross-attribute interpolation for debiasing and the potential of self-interaction strategies, which, although less stable, remain a promising direction for future exploration.
@@ 5.6. Mitigating Distribution Shift in Question Selection
While previous results have demonstrated the effectiveness of our method in improving Question Selection Module performance under OOD conditions, we further investigate the underlying mechanism by analyzing the distributional differences introduced by our approach. Specifically, Figure 6 illustrates the distribution of correct-response ratios among examinees under both the ERM baseline and our proposed method. For examinees with a lower overall correct-response ratio, our method significantly shifts the distribution toward a more moderate range, concentrating around 0.4. This suggests that fewer overly difficult questions are being selected, thereby reducing over-challenging effects. Conversely, for high-performing examinees, our method attenuates the skewness toward extremely easy questions by centering the distribution around 0.6, rather than approaching 1.0 as observed in ERM. This indicates a reduction in overfitting to bias-aligned patterns.
These results show that our method effectively mitigates the distribution shift between selected questions and the meta set, leading to a more balanced and representative question selection across diverse examinees. By increasing the diversity of bias-conflicting samples, our method adjusts the selected-question distribution to better preserve individual examinee characteristics while simultaneously aligning it more closely with a fair target distribution. This ensures that question selection reflects true proficiency rather than reinforcing spurious patterns from imbalanced training data. Our approach enhances generalization and supports more reliable proficiency estimation.
Figure 6. Distribution of Selected-Question Correct Ratios Across Examinees.
@@ 6. Conclusion
In this study, we identify and address the selection bias problem in data-driven Computerized Adaptive Testing, particularly the bias arising from meta-set distribution imbalance during training. We propose a mixup strategy guided by moderately performing students to augment underrepresented interactions and improve the generalization of the Question Selection Module. This method is simple, annotation-free, and can be seamlessly integrated into standard CAT training pipelines. Experiments on two benchmark datasets show that our approach significantly enhances both fairness and accuracy. By enhancing the diversity of bias-conflicting samples through synthetic data generation, the model learns more robust and less biased decision boundaries.This work highlights the importance of fairness in intelligent educational systems and suggests correcting bias directly from the selection process as a promising future direction.
@@ 7. GenAI Usage Disclosure
This paper involved the use of generative AI tools (e.g., ChatGPT) for non-substantive writing assistance. Specifically, AI was used to help with language polishing, improving clarity and grammar in some parts of the paper. In addition, AI-assisted tools were employed to generate code for visualizing experimental results, as well as to provide suggestions on the overall writing structure and organization. All technical content, experimental design, and core contributions were conceived and implemented by the authors.
@@ Acknowledgements.
This research was partially supported by grants from National Science and Technology Major Project under Grant (No. 2021ZD0111802), the National Natural Science Foundation of China (No. 62376086, 62406096), and the Fundamental Research Funds for the Central Universities of China (No. JZ2025HGTG0289). the China Postdoctoral Science Foundation (under Grant 2024M760722). The computation is completed on the HPC Platform of Hefei University of Technology.

### v2

@@ 5.2. Performance in OOD Settings
Motivation. In most existing data-driven CAT studies, the meta set used to evaluate question selection performance is directly sampled from each examinee’s original interaction log. This approach implicitly assumes that the observed response distribution faithfully reflects the examinee’s true exposure and proficiency profile. In practice, however, this assumption is problematic. Because CAT adaptively selects questions based on estimated proficiency, examinees are only exposed to a limited and biased subset of the entire question pool. As a result, the meta set often captures selection-induced artifacts rather than true proficiency, leading to biased and unrepresentative evaluation.
To evaluate the effectiveness of our debiasing strategy, we design an Out-of-Distribution (OOD) test setting. The training and validation splits remain identical to the IID setup; only the test partition is modified to construct the OOD scenario. Specifically, our protocol enforces a 1:1 balance of correct and incorrect responses in the meta set for each examinee, with the remaining interactions assigned to the support set. Examinees who cannot satisfy this balancing constraint are excluded. This design preserves the standard bi-level framework while deliberately introducing a distribution shift. As a result, it provides a more rigorous and interpretable benchmark for assessing model robustness and generalization, especially when evaluation is conducted under biased conditions.
As shown in Table 2, our method demonstrates clear advantages in mitigating selection bias across both datasets and diagnostic settings. When the number of selected questions increases from 5 to 10, the bias becomes more pronounced. The Worst metric drops substantially while the Avg. metric remains relatively stable, indicating that longer testing horizons disproportionately disadvantage underrepresented groups. The improvements of our approach are concentrated on the Worst metric, with relative gains ranging from 2.7% to 24.64%, whereas the Avg. metric increases by only about 0.1%–1.05%. For instance, on ASSIST0910 with NCDM-BOBCAT at Metrics@5, our method improves the Worst from 0.3662 (ERM) to 0.4989; on ASSIST0910 with IRT-BOBCAT at Metrics@10, it raises the Worst from 0.4088 (ERM) to 0.4489. For the Avg. metric, the largest gain is observed on ASSIST0910 at Metrics@10, where our method improves from 0.6091 to 0.6155, achieving a relative improvement of 1.05%. Overall, these results show that our method primarily enhances robustness by improving the performance of disadvantaged groups, while maintaining competitive overall accuracy. This aligns with the primary design goal of our debiasing framework, which is to reduce systematic disparities across different groups. At the same time, it ensures that overall performance is not sacrificed, maintaining the effectiveness of the model.
@@ 5.3. Performance in IID Settings
As shown in Table 2 and Table 3, debiasing inevitably introduces a trade-off between IID and OOD performance. While methods such as Reweight and our approach effectively mitigate bias and improve OOD generalization, they cannot surpass ERM or GroupDRO under IID settings, which are naturally more favorable to models optimized without debiasing constraints. For example, IRM achieves the best average performance on ASSIST0910 (0.6939), whereas ERM attains the highest average on NIPS-EDU (0.6930). GroupDRO also yields relatively strong IID results across both datasets. By contrast, our method prioritizes robustness: it consistently outperforms the Reweight baseline in both Worst and Avg. accuracy under OOD conditions, while remaining competitive in-distribution. These results confirm that debiasing methods, including ours, trade a small amount of IID accuracy for substantially stronger OOD generalization. This finding clearly underscores the inherent balance and trade-offs involved in robustness-oriented learning.
@@ 5.4. Hyperparameter Analysis
We conduct a hyperparameter analysis of the Mixup regularization weight $\omega$ on NIPS-EDU dataset, examining its effect on both Worst and Avg. accuracy across two Cognitive Diagnosis Models and two CAT frameworks. As shown in Figure 4, different CDM–CAT combinations exhibit distinct behaviors: NCDM-BOBCAT and IRT-UATS display a steady upward trend as $\omega$ increases, indicating that stronger regularization consistently benefits these settings, whereas IRT-BOBCAT and NCDM-UATS follow a rise-then-fall pattern, with performance improving up to a moderate $\omega$ but deteriorating when regularization becomes excessive. This suggests that while Mixup is generally effective, overly strong interpolation may introduce noise that counteracts its debiasing effect. Overall, the curves remain relatively stable across a wide range of $\omega$ values, showing that our method is not highly sensitive to this hyperparameter. Notably, the Avg. and Worst metrics do not always peak at the same $\omega$, revealing a trade-off between fairness and generalization that necessitates careful calibration to achieve balanced improvements.
Figure 4. Hyperparameter sensitivity of Mixup regularization weight $\omega$ on NIPS-EDU dataset.
$5.$ Ablation Study
We conduct ablation studies to evaluate three Mixup strategies: Mixup-Self, Mixup-Inner, and Mixup-B. In Mixup-Self, interactions are interpolated within the same examinee, introducing synthetic variation into an individual’s profile and occasionally improving robustness. Mixup-Inner interpolates the most similar examinees within the same attribute, smoothing local representations and reducing overfitting to individual-specific patterns. In contrast, Mixup-B interpolates biased examinees (e.g., from attribute A or C) with balanced counterparts (attribute B). By explicitly generating bias-conflicting interactions, this strategy enriches underrepresented patterns. It also provides a more effective and principled way to mitigate selection bias, thereby improving model robustness.
As shown in Figure 5, Mixup-B consistently achieves the best performance, particularly on the large-scale ASSIST0910 dataset, where the larger question pool provides greater diversity and amplifies the benefits of cross-attribute mixing. On NIPS-EDU, although the dataset is smaller, the same trend persists, with Mixup-B delivering the most stable improvements on both Worst and Avg.. These results demonstrate that the advantages of cross-attribute interpolation are not limited to a single dataset but generalize well across different CAT settings. At the same time, the comparison shows that Mixup-Self and Mixup-Inner can yield localized gains, but their instability makes them less reliable. Overall, the findings confirm that bias-conflicting augmentation is central to improving robustness and suggest promising directions for hybrid designs that combine multiple Mixup strategies in future work.
@@ 5.6. Mitigating Distribution Shift in Question Selection
While previous results demonstrate the effectiveness of our method in improving Question Selection Module performance under OOD conditions, we further investigate its mechanism by analyzing the distributional differences it introduces. Specifically, Figure 6 presents the distribution of correct-response ratios among examinees under both the ERM baseline and our method. For examinees in attribute A, our method shifts the distribution toward a more moderate range (around 0.4), indicating that fewer overly difficult questions are selected and the risk of over-challenging is reduced. For examinees in attribute C, the skewness toward extremely easy questions is effectively alleviated, with the distribution centered around 0.6 instead of approaching 1.0 as in ERM, suggesting less overfitting to bias-aligned patterns overall.
Overall, these results demonstrate that our method effectively reduces the gap between the selected-question distribution and the meta set, yielding a more balanced and representative selection across examinees. By enriching bias-conflicting samples, it better preserves individual characteristics while aligning selection with a fairer target distribution, ensuring that proficiency estimation reflects true ability rather than artifacts of imbalanced training data. In turn, this leads to improved generalization, greater robustness, and more reliable assessment outcomes.
Figure 6. Distribution of selected-question correct-response ratios across examinees.
@@ 6. Conclusion
In this study, we identify and address the problem of selection bias in data-driven Computerized Adaptive Testing, focusing on the imbalance of meta-set distributions during training. We propose a Mixup strategy guided by moderately performing examinees to augment underrepresented interactions and enhance the generalization of the Question Selection Module. The method is simple, effective, and can be seamlessly integrated into data-driven CAT training pipelines. Experiments on two benchmark datasets demonstrate that our approach mitigates bias while improving accuracy. By enriching bias-conflicting samples through synthetic data generation, the model learns more robust and less biased decision boundaries. This work highlights the importance of bias mitigation in intelligent educational systems and suggests addressing bias directly within the selection process as a promising direction for future research. Further investigations may explore integrating debiasing strategies with other CAT components to achieve broader improvements in reliability and validity.
@@ 7. GenAI Usage Disclosure
This paper used generative AI tools (e.g., ChatGPT) for non-substantive assistance such as language polishing, grammar correction, simple code generation, and writing suggestions. All technical content, experiments, and core contributions were conceived and implemented solely by the authors.
@@ Acknowledgements.
This research was partially supported by grants from the National Science and Technology Major Project (No. 2021ZD0111802), the National Natural Science Foundation of China (Nos. 62376086, 62406096), the Fundamental Research Funds for the Central Universities of China (No. JZ2025HGTG0289), and the China Postdoctoral Science Foundation (No. 2024M760722). The computations were completed on the HPC Platform of Hefei University of Technology.